Jim and Wanda both have some apples. If Jim gives Wanda an apple, they will both have the same number of apples. However, if Wanda gives Jim an apple, Jim will have twice as many as Wanda. How many apples does Jim have and how many apples does Wanda have?
Answer by stanbon(75887)

Jim and Wanda both have some apples. If Jim gives Wanda an apple, they will both have the same number of apples. However, if Wanda gives Jim an apple, Jim will have twice as many as Wanda. How many apples does Jim have and how many apples does Wanda have?
-------
Equations:
j-1 = w+1
j = 2(w-1)
------
Substitute for "j" and solve for "w":
2(w-1) - 1 = w + 1
2w-2 -1 = w+1
2w-3 = w+1
w = 4 (Wanda has 4 apples)
----
solve for "j":
j = 2(w-1)
j = 2(3) = 6 (Jim has 6 apples).
---------------------------
Cheers,
Stan H.
=============